OpenStudy (anonymous):

find the volume bounded by x^2 + y^2 + z^2 = 4 and r = cos theta and r = sin theta

OpenStudy (anonymous):

\[ {x}^2 + {y}^2 + {z}^2 = 4 r = \cos \theta r = \sin \theta\]

OpenStudy (anonymous):

there's supposed to be spaces between each function.

OpenStudy (anonymous):

Please help!